[figure: 9.7 cm]

ELECTRON TRANSPORT AND POINT CONTACT SPECTROSCOPY OF CUBIC DyCu_{5}

K. Nenkov^{a,b}, B. Idzikowski^{c}, S. Ilkovič^{d}, E. Kačmarčiková^{e}, M. Reiffers^{e} and K.-H. Müller^{a} aInstitut für Festkörper- und Werkstofforschung Dresden, Germany

bInternational Laboratory of High Magnetic Fields and Low Temp., Wrocław, Poland

cInstitute of Molecular Physics, Polish Academy of Sciences, Poznań, Poland dFaculty of Sciences, University of Prešov, Prešov, Slovakia

eInstitute of Experimental Physics, Slovak Academy of Sciences, Košice, Slovakia

The electron transport properties and point contact spectra of the cubic phase (structure type AuBe_{5}, space group F-43m) of melt-spun DyCu_{5} have been investigated. In acsusceptibility and magnetization measurements, it has been found that below a transition temperature T_{m}=6.5 K this phase is metamagnetic. The samples show a spontaneous magnetization between T_{m} and T_{C}=14.5 K. These magnetic phenomena imply also uncommon electron transport effects. Our measurements of electrical resistivity confirm the existence of two magnetic phase transitions in the compound. This magnetic behavior is attributed to RKKY-type indirect exchange interactions. Further, we present point contact spectra of DyCu_{5} - Cu heterocontacts which are directly proportional to the electron-quasiparticle interaction function. Apart of a peak at about 18 meV, which is characteristic for electron-phonon interaction in pure Cu, we have observed a series of additional maxima at lower energy, which are connected with the allowed 4f-transitions from the ground state to excited crystalline electric field levels.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : K. Nenkov

Address for correspondence : Institut für Festkörper- und Werkstofforschung Dresden, Postfach 270016, 01171 Dresden, Germany

Email address : Staiger@ifw-dresden.d400.de